

********************************************************************
UARS WEEKLY STATUS REPORT - 3 Dec 2004
********************************************************************
***********
SOLSTICE
***********
SOLSTICE began day only operations on Dec 2nd in support of the yaw
maneuver. Otherwise, the instrument continues to perform exceptionally,
taking solar data on all available orbital tracking periods. An average
of 57% of data were returned this week accounting for 364 minutes of
solar observing time.
***********
SUSIM
***********
SUSIM was powered down on December 1 for the yaw-around.
Beforehand, the experiment successfully gathered daily mid (1.1 nm)
resolution, MgII, occultation, solar pointing offset, and continuous
monitoring scans for time periods with UARS telemetry or tape
recording.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
The HALOE Project at NASA Langley met on December 1 and
discussed activities related to (1) its ongoing assessment of
possible instrument trends that could affect its retrieved results;
(2) ACE/HALOE comparisons obtained during mid 2004 at
high northern latitudes; and (3) activities related to its continuing
operations and the processing of its measurement events. Ellis
Remsberg visited with UARS Project personnel at GSFC on
November 19. Jim Russell and John Anderson of Hampton University
and Ellis Remsberg of NASA Langley will be making presentations
at the Fall AGU meeting. HALOE Team members also discussed
the science priorities for HALOE in preparation for the UARS
response to NASA's upcoming extended mission opportunity.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 20 November 2004 (Orbit 72203, GMT Day 325) through
26 November 2004 (Orbit 72308, GMT Day 331).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
-NONE THIS REPORT PERIOD.
The beta angle ranged between the angles of 35.2 degrees to 21.7
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 02 December 2004 (DOY 337).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.27 to 1.29 degrees C. Additional
battery performance data for each battery on 19 November and 26
November 2004 are provided in Appendix D.
The clock error ranged from 10.0 to -12.0 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 02
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continued to work on integrating the SMA
functionality. More testing was done regarding the file transfer
issue between the User Planning System and CMS to help isolate the
problem. Testing this week showed that the USM file is being
manipulated before it gets to the CMS from the UPS. We believe this
to be an issue with the FORMATS system. The FOT will call a meeting
early next week to get FORMATS' software development involved.
MEETINGS
The FOT called a meeting with CMS software development and the CMS
system administrator on 24 November 2004 to discuss the SMA problems.
At this meeting the status of the investigation was reviewed and
future tests and plans were outlined.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 02 December 2004 (DOY 337).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 26 November 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
72209 325/1034 04-325 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
72227 326/1437 04-326 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
72293 331/0049 04-327 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 26 November 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 26 November 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
72209 325/1036 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
72227 326/1506 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
72293 331/0054 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 26 November 2004
19 November 2004 - GMT Day 324 Beta = 37.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.42/-3.68 +00.0/+00.0 07.0 0.00 0.0
2 -05.6/ -95.2 +6.43/+5.46 +27.2/-15.2 26.2 1.04 19.7
3 -11.2/-168.0 +5.14/+4.50 +25.6/-14.4 26.2 1.03 19.0
26 November 2004 - GMT Day 331 Beta = 21.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.74/-3.99 +00.0/+00.0 07.0 0.00 0.0
2 22.4/ -50.4 +6.43/+5.14 +28.8/-14.4 26.2 1.04 19.3
3 11.2/ -67.2 +5.14/+3.87 +28.0/-14.8 26.1 1.04 19.0
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 26 November 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-086 FOT 72254 328/0914 30832 Generic Late Acquisition
04-089 FOT 72289 330/1816 37679 TDRSS-E Data Dropout
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-088 FOT 72109 318/1856 N/A File Transfer Problem
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-086 FOT See List Nov 04 30832 Generic Late Acquisition
04-089 FOT 72289 330/1816 37679 TDRSS-E Data Dropout